OpenStudy (anonymous):

how do I convert y=2x^2+2x-24 into vertex form?

OpenStudy (anonymous):

Y= 2(x+1/4)^2 -49/2

OpenStudy (anonymous):

isn't vertex form y-k= a(x-h)^2

OpenStudy (anonymous):

its y=a(x-h)^2+k

jimthompson5910 (jim_thompson5910):

both forms are essentially equivalent because you can add k to both sides to go from y-k = a(x-h)^2 to y = a(x-h)^2+k

OpenStudy (anonymous):

Ok thank you!! (: